# On Mean Field Games in Infinite Dimension

### v2

@@ 1. Introduction
(HJB) $-{\partial_{t}}v(t,x)-{L}v(t,x)+\mathcal{H}(x,Dv(t,x),m(t))=0,\ \ \ v(T,\cdot)=G(\cdot,m(T)),$
(FP) $\partial_{t}m(t)-{L}^{*}m(t)-\mbox{div}\Big{(}\mathcal{H}_{p}(x,Dv(t,x),m(t))\,m(t)\Big{)}=0,\ \ \ m(0)=m_{0},$
@@ 2. Preliminaries and Assumptions
[…]
@@ 2.3. The operator $A$ and the generated semigroup
    It can be verified from (2.7) that Assumption 2.1 implies typical conditions on the operator $A$ ensuring existence and uniqueness of mild solutions of equations in infinite dimensional spaces (see [21, Ch. 4]), that is:
@@ 2.4. The probabilistic framework and the reference SDE
Under Assumption 2.1, the stochastic convolution (see Remark 2.2(ii) and [20, Ch. 5])
@@
(2.10) $[R_{t}\varphi](x)=\mathsf{E}\left[\varphi(X_{t}^{0,x})\right]=\int_{H}\varphi(y)\mathcal{N}(e^{tA}x,Q_{t})(\mathrm{d}y),\ \ \ \varphi\in C_{b}(H),$
called the Ornstein-Uhlenbeck semigroup. Above, $\mathcal{N}(a,Q)$ is the Gaussian measure in $H$ with mean $a\in H$ and covariance operator $Q\in L^{1}_{+}(H)$ (see, e.g., [19, Ch.2, Sec.3]). We will use it later to define the concept of a mild solution to (HJB).
@@
When $w$ does not depend on $t$, the result is proved [16, Th. 13] and the proof works in the same way if $w$ depends on $t$. We just need to explain how our setup fits into the framework of [16].
Given our filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathsf{P})$ with the Wiener process $W$, let $(\Omega_{1},\mathcal{F}_{1},\mathsf{P}_{1})$ be another complete atomless probability space. For any any $m_{0}\in\mathcal{P}_{1}(H)$, since $(\Omega_{1},\mathcal{F}_{1},\mathsf{P}_{1})$ is nonatomic, there exists an $\mathcal{F}_{1}$-measurable random variable $\xi_{0}:\Omega_{1}\rightarrow X$ such that $m_{0}=\mathcal{L}(\xi_{0})$ (see, e.g., [11, Lemma 5.29]). We define a new probability space $(\Omega_{2},\mathcal{F}_{2},\mathsf{P}_{2})$ as follows:
@@
Thus, the law of $X_{0}$ in the new probability space is the same. This puts us in the framework of [16] and we can now proceed as in the proof of [16, Th. 13].
@@
    Using the definition of mild solution, the fact that $|e^{tA}|\leq 1$ and the estimate on the stochastic convolution $W_{t}^{A}$ [21, Prop. 1.144], we have, for some $c_{1}>0$,
@@
[…]
@@
      The map $[0,T]\rightarrow L^{2}(\Omega,\mathsf{P})$, $t\mapsto W^{A}_{t}$ is continuous (see [20, Th. 5.2]); hence, also the map $[0,T]\rightarrow L^{1}(\Omega,\mathsf{P})$, $t\mapsto W^{A}_{t}$ is continuous. Hence, since $[0,T]$ is compact, there exists a modulus of continuity $\hat{\omega}_{A}$ such that, for every $0\leq s\leq t\leq T$,
@@ 3. The Fokker-Planck Equation: Weak Solutions
for every $t\in[0,T]$ and $\varphi\in\mathcal{D}_{T}$. [^1] Instead of the requirement that $m\in\mathcal{S}$, in [6] it was assumed that a weak solution $m:[0,T]\rightarrow\mathcal{P}_{1}(H)$ is such that $t\mapsto m(t)(I)$ is measurable for each $I\in\mathcal{B}(H)$, that is, in the terminology of [6], $m$ is a probability kernel.
@@
This is just an application of Dynkin’s formula for sufficiently regular functions: see, e.g., [21, Prop. 1.169], applied there with
$b_{0}(s,x,a_{1})=a_{1},\ a_{1}(s,\omega)=w(s,X^{w}_{s}(\omega)),\ a_{2}(\cdot)=0,\ \Lambda=B_{\|w\|_{\infty}}\subset H.\vspace{-.9cm}$
@@
The claim follows from [6, Th. 4.1] once we prove that a solution in our sense is also a solution in the sense of [6, Th. 4.1]. To do that we first notice that, in [6, Formula (1.5)], the definition of solution is given using the set test functions (called $\mathcal{E}_{A}(H)$ there) which is the linear span of the real parts of the functions of the form
@@ 4. The HJB Equation: Mild Solutions
In this section we deal with equation (HJB) for a given fixed $m\in\mathcal{S}$. We introduce the concept of a mild solution to (HJB) (see [21, Section 4.4.1.2]).
@@
    This part is contained in [21, Theorem 4.90]. [^2] More precisely, it is enough to take in that theorem, $m=0$, $\gamma_{G}(t)=t^{-\gamma}$, $G=\mathbf{Id}$. We warn that $G$ in this footnote refers to the operator used in [21, Theorem 4.90], which has nothing to do with our function $G$.
@@
    Using the smoothing properties of the Ornstein-Uhlenbeck semigroup (see e.g. [21, Theorem 4.37]), we have the representation of the gradients
@@
    Hence, fixing $t\in[0,T)$ and considering [21, Th. 4.37] and Assumption 2.6, we have, for some $C>0$,
@@
    By a suitable generalization of Gronwall’s Lemma (see [21, Prop. D.30]), we get
@@ 5. The MFG System: Existence of Solutions
Observe that, by Lemma A.1 and by [41, Remark 7.13 (ii)], the metric $\mathbf{d}_{1}$ metrizes on $\mathcal{R}_{m_{0}}$ the topology of weak convergence in $\mathcal{P}(H)$. In view of Proposition 5.2, we also define
@@
We will use the fact (see, e.g., [41, Rem. 7.5(i)] or [11, Cor. 5.4]) that for $\mu,\nu\in\mathcal{P}_{1}(H)$
@@
where ${\rm Lip}_{1}$ is the set of Lipschitz functions on $H$ with Lipschitz constant $1$. Then
@@
    Since $\mathbf{d}_{1}$ metrizes the weak topology in $\mathcal{Q}_{m_{0}}$, for each $i=1,...,N$ we have the existence of $\delta>0$ and a family $\{f_{i,1},...,f_{i,M}\}\subset C_{b}(H)$ such that
    $\mathcal{V}_{t_{i}}(\bar{m}(t_{i}))$ :=$:=\{\mu\in\mathcal{Q}_{m_{0}}:\ |\mu-\bar{m}(t_{i})|_{f_{i,j}}<\delta,\ \ \ \forall j=1,...,M\}$
@@
    $\mathcal{V}_{N}(\bar{m}(\cdot))$ :=$:=\Big{\{}m(\cdot)\in\mathcal{C}_{m_{0}}:\ |m(t_{j})-\bar{m}(t_{j})|_{f_{i}}<\varepsilon/2,\ \ \forall i=1,...,N,\ \forall j=1,...,M\Big{\}}$
@@
    Since $\mathbf{d}_{1}$ metrizes the weak topology on $\mathcal{Q}_{m_{0}}$, for each $j=1,...,M$ we have the existence of $\delta_{j}>0$ such that
@@
    $\mathcal{U}(\bar{m}(\cdot))$ :=$:=\left\{m(\cdot)\in\mathcal{C}_{m_{0}}:\ \sup_{t\in[0,T]}\mathbf{d}_{1}(m(t),\bar{m}(t))<\delta\right\}$
@@ 6. The MFG System: Uniqueness of Solutions
In this section, we establish the uniqueness of solutions for our MFG system (HJB)–(FP). It is well known, even in finite-dimensional state spaces (see [10, Rem. 1.4, p.32]), that uniqueness of a solution for this system cannot generally be guaranteed. A typical approach to ensure uniqueness is to impose certain monotonicity conditions on the data. One of the most common monotonicity conditions is the Lasry-Lions monotonicity condition (see, for example, [33, 34, 35] and [10, Section 1.3.2]). For other monotonicity conditions that ensure uniqueness in finite-dimensional MFG systems, readers can refer to [31, 37] and the references therein. Another typical feature of uniqueness results is the separability of the Hamiltonian function $\mathcal{H}$ (as assumed in Assumption 6.1(i) here). Well-posedness of MFG systems without separability has been studied in [37]. Here, we assume separability and use a variant of the Lasry-Lions monotonicity condition from [10, Theorem 1.4]. Our assumptions are as follows.
@@
Uniqueness. The proof is an adaptation to our Hilbert space setting of a typical proof of uniqueness in finite dimensional spaces (see, e.g., [10, Theorem 1.4]).
@@
[…]
@@
$+\int_{0}^{T}\int_{H}\left[\mathcal{H}^{0}(x,Dv_{2}(s,x))-\mathcal{H}^{0}(x,Dv_{1}(s,x))-\langle\mathcal{H}^{0}_{p}(x,Dv_{1}(s,x)),-D\bar{v}(s,x)\rangle\right]m_{1}(s,\mathrm{d}x)$
$+\int_{0}^{T}\int_{H}\left[\mathcal{H}^{0}(x,Dv_{1}(s,x))-\mathcal{H}^{0}(x,Dv_{2}(s,x))-\langle\mathcal{H}^{0}_{p}(x,Dv_{2}(s,x)),\,D\bar{v}(s,x)\rangle\right]m_{2}(s,\mathrm{d}x).$
By Assumption 6.1(ii)-(iii) all terms above are nonnegative, so they all must be equal to 00. We now conclude as follows.
@@ 7. Examples from Stochastic Optimal Control
[…]
@@
Let $\ell:H\times[0,\infty)\rightarrow\mathbb{R}$ be bounded, continuous, and such that $\ell(x,\cdot)$ is strictly increasing for every $x\in H$ and is Lipschitz continuous with Lipschitz constant independent of $x$. Let $\rho:H\rightarrow[0,\infty)$ be bounded and Lipschitz continuous. Let $\nu$ be a positive finite measure on $H$ which is full  [^5] That is $\nu(D)>0$ for every open set $D$ in $H$. An example of a measure which is full is the nondegenerate Gaussian measure with mean $a\in H$ and covariance operator $Q\in\mathcal{L}_{1}^{+}(H)$, see [15, Proposition 1.25] (nondegeneracy means that ${\rm ker}\,Q=\{0\}$).. We define
@@
because $\ell(z,\cdot)$ is increasing. Moreover, if the last expression is equal to 00, using that $\ell(z,\cdot)$ is strictly increasing, we conclude that $\rho*\mu_{1}(z)-\rho*\mu_{2}(z)=0,\mu$ a.e.. However, since this function is continuous and $\mu$ is full, this implies that $\rho*\mu_{1}(z)=\rho*\mu_{2}(z)$ for every $z\in H$. Hence, $F(\cdot,\mu_{1})=F(\cdot,\mu_{2})$ on $H$. Therefore, $F$ satisfies Assumption 6.1(iii) and Assumption 6.1(iv)(b). Finally, we compute

### v3

@@ 1. Introduction
(HJB) $-{\partial_{t}}v(t,x)-{L}v(t,x)+\mathcal{H}(x,Dv(t,x),m(t))=0,\ \ \ v(T,\cdot)=G(\cdot,m(T)),\ \ \ \ \$
(FP) $\partial_{t}m(t)-{L}^{*}m(t)-\mbox{div}\Big{(}\mathcal{H}_{p}(x,Dv(t,x),m(t))\,m(t)\Big{)}=0,\ \ \ m(0)=m_{0},\ \ \ \ \ \$
@@ 2. Preliminaries and Assumptions
We recall that $H$ is a real separable Hilbert space with inner product $\langle\cdot,\cdot\rangle$. We will write $|\cdot|$ for the norm in $H$; if needed we stress $H$ in the notation writing $|\cdot|_{H}$. We always identify $H$ with its dual space. For $R>0$, we denote $B_{R}:=\{x\in H:|x|\leq R\}$. Throughout the paper, $\mathbb{N}$ will denote the set of natural numbers without 0.
@@ 2.3. The operator $A$ and the generated semigroup
    It can be verified from (2.7) that Assumption 2.1 implies typical conditions on the operator AA ensuring existence and uniqueness of mild solutions of equations in infinite dimensional spaces (see [22, Ch. 4]), that is:
@@ 2.4. The probabilistic framework and the reference SDE
Under Assumption 2.1, the stochastic convolution (see Remark 2.2(ii) and [21, Ch. 5])
@@
(2.10) $R_{t}\varphi](x)=\mathsf{E}\left[\varphi(X_{t}^{0,x})\right]=\int_{H}\varphi(y)\mathcal{N}(e^{tA}x,Q_{t})(\mathrm{d}y),\ \ \ \varphi\in C_{b}(H),$
called the Ornstein-Uhlenbeck semigroup. Above, $\mathcal{N}(a,Q)$ is the Gaussian measure in $H$ with mean $a\in H$ and covariance operator $Q\in L^{1}_{+}(H)$ (see, e.g., [20, Ch.2, Sec.3]). We will use it later to define the concept of a mild solution to (HJB).
@@
When $w$ does not depend on $t$, the result is proved [17, Th. 13] and the proof works in the same way if $w$ depends on $t$. We just need to explain how our setup fits into the framework of [17].
Given our filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathsf{P})$ with the Wiener process $W$, let $(\Omega_{1},\mathcal{F}_{1},\mathsf{P}_{1})$ be another complete atomless probability space. For any any $m_{0}\in\mathcal{P}_{1}(H)$, since $(\Omega_{1},\mathcal{F}_{1},\mathsf{P}_{1})$ is nonatomic, there exists an $\mathcal{F}_{1}$-measurable random variable $\xi_{0}:\Omega_{1}\rightarrow X$ such that $m_{0}=\mathcal{L}(\xi_{0})$ (see, e.g., [12, Lemma 5.29]). We define a new probability space $(\Omega_{2},\mathcal{F}_{2},\mathsf{P}_{2})$ as follows:
@@
Thus, the law of $X_{0}$ in the new probability space is the same. This puts us in the framework of [17] and we can now proceed as in the proof of [17, Th. 13].
@@
    Using the definition of mild solution, the fact that $|e^{tA}|\leq 1$ and the estimate on the stochastic convolution $W_{t}^{A}$ [22, Prop. 1.144], we have, for some $c_{1}>0$,
@@
      Let $R>0$. Since $A$ generates a semigroup of contractions, for every $0<\eta<s\leq T$ and $w$ such that $|w|_{\infty}\leq R$, we have ($\mathsf{P}\mbox{-a.s.}$)
@@
      The map $0,T]\rightarrow L^{2}(\Omega,\mathsf{P})$, $t\mapsto W^{A}_{t}$ is continuous (see [21, Th. 5.2]); hence, also the map $0,T]\rightarrow L^{1}(\Omega,\mathsf{P})$, $t\mapsto W^{A}_{t}$ is continuous. Hence, since $[0,T]$ is compact, there exists a modulus of continuity $\hat{\omega}_{A}$ such that, for every $0\leq s\leq t\leq T$,
@@ 3. The Fokker-Planck Equation: Weak Solutions
for every $t\in[0,T]$ and $\varphi\in\mathcal{D}_{T}$. [^1] Instead of the requirement that $m\in\mathcal{S}$, in [7] the notion of weak solution $m:[0,T]\rightarrow\mathcal{P}_{1}(H)$ requires that $t\mapsto m(t)(I)$ is measurable for each $I\in\mathcal{B}(H)$; that is, in the terminology of [7], $m$ is a probability kernel. Our requirement that $m\in\mathcal{S}$ is actually stronger. Indeed, given $I\in\mathcal{B}(H)$ and $m\in\mathcal{S}$, the map $t\mapsto m(t,I)$ is the composition of the continuous map $0,T]\rightarrow(\mathcal{P}_{1}(H),\mathbf{d}_{1}),\ t\mapsto m(t)$ and of the map $(\mathcal{P}_{1}(H),\mathbf{d}_{1})\rightarrow\mathbb{R},\ \mu\mapsto\mu(I)$. The $\mathbf{d}_{1}$-topology is stronger than the weak topology of measures; so, to conclude that $m\in\mathcal{S}$ is a probability kernel, it suffices to show that $\mu\mapsto\mu(I)$ is measurable when $\mathcal{P}_{1}(H)$ is endowed with the Borel $\sigma$-algebra induced by the weak topology. This claim follows from the fact that the latter map is upper semicontinuous for $I$ closed (that is standard, as a consequence of one of the equivalent definitions of weak convergence) and by [4, Prop. 7.25]. This makes our notion of solution stronger and enables us to invoke the uniqueness result of [7] in Proposition 3.3.
@@
This is just an application of Dynkin’s formula for sufficiently regular functions: see, e.g., [22, Prop. 1.169], applied there with
$b_{0}(s,x,a_{1})=a_{1},\ a_{1}(s,\omega)=w(s,X^{w}_{s}(\omega)),\ a_{2}(\cdot)=0,\ \Lambda=B_{\|w\|_{\infty}}\subset H.\vskip-25.6073pt$
@@
The claim follows from [7, Th. 4.1] once we prove that a solution in our sense is also a solution in the sense of [7, Th. 4.1]. To do that we first notice that, in [7, Formula (1.5)], the definition of solution is given using the set test functions (called $\mathcal{E}_{A}(H)$ there) which is the linear span of the real parts of the functions of the form
@@ 4. The HJB Equation: Mild Solutions
In this section we deal with equation (HJB) for a given fixed $m\in\mathcal{S}$. We introduce the concept of a mild solution to (HJB) (see [22, Section 4.4.1.2]).
@@
    This part is contained in [22, Theorem 4.90]. [^2] More precisely, it is enough to take in that theorem, $m=0$, $\gamma_{G}(t)=t^{-\gamma}$, $G=\mathbf{Id}$. We warn that $G$ in this footnote refers to the operator used in [22, Theorem 4.90], which has nothing to do with our function $G$.
@@
    Using the smoothing properties of the Ornstein-Uhlenbeck semigroup (see e.g. [22, Theorem 4.37]), we have the representation of the gradients
@@
    Hence, fixing $t\in[0,T)$ and considering [22, Th. 4.37] and Assumption 2.6, we have, for some $C>0$,
@@
    By a suitable generalization of Gronwall’s Lemma (see [22, Prop. D.30]), we get
@@ 5. The MFG System: Existence of Solutions
Observe that, by Lemma A.1 and by [42, Remark 7.13 (ii)], the metric $\mathbf{d}_{1}$ metrizes on $\mathcal{R}_{m_{0}}$ the topology of weak convergence in $\mathcal{P}(H)$. In view of Proposition 5.2, we also define
@@
We will use the fact (see, e.g., [42, Rem. 7.5(i)] or [12, Cor. 5.4]) that for $\mu,\nu\in\mathcal{P}_{1}(H)$
@@
where ${\rm Lip}_{1}$ is the set of Lipschitz functions on $H$ with Lipschitz constant 11. Then
@@
    Since $\mathbf{d}_{1}$ metrizes the weak topology in $\mathcal{Q}_{m_{0}}$, for each i=1,,Ni=1,...,N we have the existence of $\delta>0$ and a family $\{f_{i,1},...,f_{i,M}\}\subset C_{b}(H)$ such that
    $\mathcal{V}_{t_{i}}(\bar{m}(t_{i}))$ $:=\{\mu\in\mathcal{Q}_{m_{0}}:\ |\mu-\bar{m}(t_{i})|_{f_{i,j}}<\delta,\ \ \ \forall j=1,...,M\}$
@@
    $\mathcal{V}_{N}(\bar{m}(\cdot))$ $:=\Big{\{}m(\cdot)\in\mathcal{C}_{m_{0}}:\ |m(t_{j})-\bar{m}(t_{j})|_{f_{i}}<\varepsilon/2,\ \ \forall i=1,...,N,\ \forall j=1,...,M\Big{\}}$
@@
    Since $\mathbf{d}_{1}$ metrizes the weak topology on $\mathcal{Q}_{m_{0}}$, for each j=1,,Mj=1,...,M we have the existence of $\delta_{j}>0$ such that
@@
    $\mathcal{U}(\bar{m}(\cdot))$ $:=\left\{m(\cdot)\in\mathcal{C}_{m_{0}}:\ \sup_{t\in[0,T]}\mathbf{d}_{1}(m(t),\bar{m}(t))<\delta\right\}$
@@ 6. The MFG System: Uniqueness of Solutions
In this section, we establish the uniqueness of solutions for our MFG system (HJB)–(FP). It is well known, even in finite-dimensional state spaces (see [11, Rem. 1.4, p.32]), that uniqueness of a solution for this system cannot generally be guaranteed. A typical approach to ensure uniqueness is to impose certain monotonicity conditions on the data. One of the most common monotonicity conditions is the Lasry-Lions monotonicity condition (see, for example, [34, 35, 36] and [11, Section 1.3.2]). For other monotonicity conditions that ensure uniqueness in finite-dimensional MFG systems, readers can refer to [32, 38] and the references therein. Another typical feature of uniqueness results is the separability of the Hamiltonian function $\mathcal{H}$ (as assumed in Assumption 6.1(i) here). Well-posedness of MFG systems without separability has been studied in [38]. Here, we assume separability and use a variant of the Lasry-Lions monotonicity condition from [11, Theorem 1.4]. Our assumptions are as follows.
@@
Uniqueness. The proof is an adaptation to our Hilbert space setting of a typical proof of uniqueness in finite dimensional spaces (see, e.g., [11, Theorem 1.4]).
@@
$\begin{cases}{\partial_{t}}{v}(t,x)+{L}{v}(t,x)=f(t,x),\vskip 6.0pt plus 2.0pt minus 2.0pt\\ v(T,x)=\phi(x):=G(x,m_{1}(T))-G(x,m_{2}(T)),\end{cases}$
@@
$\int_{0}^{T}\int_{H}\left[\mathcal{H}^{0}(x,Dv_{2}(s,x))-\mathcal{H}^{0}(x,Dv_{1}(s,x))-\langle\mathcal{H}^{0}_{p}(x,Dv_{1}(s,x)),-D\bar{v}(s,x)\rangle\right]m_{1}(s,\mathrm{d}x)$
$\int_{0}^{T}\int_{H}\left[\mathcal{H}^{0}(x,Dv_{1}(s,x))-\mathcal{H}^{0}(x,Dv_{2}(s,x))-\langle\mathcal{H}^{0}_{p}(x,Dv_{2}(s,x)),\,D\bar{v}(s,x)\rangle\right]m_{2}(s,\mathrm{d}x).$
By Assumption 6.1(ii)-(iii) all terms above are nonnegative, so they all must be equal to 0. We now conclude as follows.
@@ 7. Examples from Stochastic Optimal Control
$\mathcal{H}^{1}(p)=\begin{cases}(f_{1}^{\prime})^{-1}(|p|)|p|-f_{1}((f_{1}^{\prime})^{-1}(|p|)),\quad\mbox{if}\,\,|p|<f_{1}^{\prime}(R),\vskip 6.0pt plus 2.0pt minus 2.0pt\\ R|p|-f_{1}(R),\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \quad\mbox{if}\,\,|p|\geq f_{1}^{\prime}(R).\end{cases}$
@@
Let $\ell:H\times[0,\infty)\rightarrow\mathbb{R}$ be bounded, continuous, and such that $\ell(x,\cdot)$ is strictly increasing for every $x\in H$ and is Lipschitz continuous with Lipschitz constant independent of $x$. Let $\rho:H\rightarrow[0,\infty)$ be bounded and Lipschitz continuous. Let $\nu$ be a positive finite measure on $H$ which is full  [^5] That is $\nu(D)>0$ for every open set $D$ in $H$. An example of a measure which is full is the nondegenerate Gaussian measure with mean $a\in H$ and covariance operator $Q\in\mathcal{L}_{1}^{+}(H)$, see [16, Proposition 1.25] (nondegeneracy means that ${\rm ker}\,Q=\{0\}$).. We define
@@
because $\ell(z,\cdot)$ is increasing. Moreover, if the last expression is equal to 0, using that $\ell(z,\cdot)$ is strictly increasing, we conclude that $\rho*\mu_{1}(z)-\rho*\mu_{2}(z)=0,\mu$ a.e.. However, since this function is continuous and $\mu$ is full, this implies that $\rho*\mu_{1}(z)=\rho*\mu_{2}(z)$ for every $z\in H$. Hence, $F(\cdot,\mu_{1})=F(\cdot,\mu_{2})$ on $H$. Therefore, $F$ satisfies Assumption 6.1(iii) and Assumption 6.1(iv)(b). Finally, we compute